A Carnot engine works between two heat reservoirs at temperatures Thot=300.0K and Tcold=230.0K.

What is its efficiency?

If it absorbs 140.0J from the hot reservoir during each cycle, how much work does it do?

How much heat does it reject during each cycle?

What is the coefficient of perfomance, COP, of this engine when it works as a refrigerator between the same two reservoirs?
Answer #1

Efficiency of carnot engine is

\eta = 1 - \left ( \frac{T_2}{T_1} \right )

\eta = 1 - \left ( \frac{230}{300} \right )

\eta = 0.233\approx 23.3\%

b) The work done in terms of heat and efficiency is

W = Q \eta = (140)0.233 = 32.67 J

c) Coefficient of performance in terms of efficiency is

K = \frac{1}{\eta} - 1 = \frac{1}{0.233} - 1 \Rightarrow 3.29